A basketballer is trying to make a shot in an international game from the 3 point line (6.75 m
away from the hoop). The basketball hoop is 3.05 m above the floor, the basketballer’s hands
are 1.95 m above the floor when they release the ball at an angle of 39° above the horizontal.
What velocity would the ball have to have been thrown with to reach and fall down into the
hoop? And if the ball has a mass of 625 g what was the ball’s maximum potential energy during
the throw?
1
Expert's answer
2019-10-17T11:15:26-0400

The equation of trajectory of the ball when it moves as projectile


"y=y_0+\tan(\theta)\cdot x-\frac{g}{2v_0^2\cos^2(\theta)}\cdot x^2"

In our case we have an equation


"3.05=1.95+\tan(39^{\circ})\cdot 6.75-\frac{9.8}{2v_0^2\cos^2(39^{\circ})}\cdot 6.75^2"

So, the initial velocity


"v_0=9.2 \:\rm m/s"

The maximum height


"y_{max}=y_0+\frac{v_0^2\sin^2(\theta)}{2g}"

"=1.95+\frac{9.2^2\sin^2(39^{\circ})}{2\times 9.8}=3.66\:\rm m"

So, the ball’s maximum potential energy during the throw


"E_{max}=mgy_{max}"

"=0.625\:\rm kg\times 9.8\: N/kg\times 3.66\: m=22.4\:\rm J"